OpenStudy (anonymous):

f(x)=1/27-3^x for which value is x undefined?

OpenStudy (bahrom7893):

it's defined everywhere

OpenStudy (anonymous):

im looking for when its undefined..

OpenStudy (bahrom7893):

well there are no parts when its undefined...

OpenStudy (anonymous):

would it be 3? because when three is plugged in the denominator = 0

OpenStudy (bahrom7893):

OH, can you write properly then?

OpenStudy (bahrom7893):

look use parenthesis, there's a difference b/n f(x)=1/27-3^x and f(x)=1/(27-3^x)

OpenStudy (bahrom7893):

then yeah it would be undefined at x=3